OpenStudy (kitkat1):

A pair of equations is shown below: y = 8x – 9 y = 4x – 1 Part A: Explain how you will solve the pair of equations by substitution or elimination. Show all the steps and write the solution.

OpenStudy (kitkat1):

I will give medals Please help me

OpenStudy (yadu123):

Set the equation equal to each other

OpenStudy (kitkat1):

how can i do that im not really good at that

OpenStudy (yadu123):

Since both of them are equal to y you set the equation to be 8x-9=4x-1

OpenStudy (kitkat1):

okay and then what?

OpenStudy (yadu123):

Combine Like terms. Bring 4x to the other side by subtracting and combine like terms

OpenStudy (kitkat1):

so minus the 4 and what other number?

OpenStudy (yadu123):

So 8x-4x=4x-9=-1

OpenStudy (yadu123):

Since 8x and 4x share a x they are like terms and you can combine them

OpenStudy (yadu123):

So now solve 4x-9=-1

OpenStudy (kitkat1):

so now we have 4x - 9 = -1

OpenStudy (yadu123):

yes. Bring nine to the other side by adding and divide by 4

OpenStudy (kitkat1):

does the 1 has to be negative?

OpenStudy (yadu123):

yes since it is negative from the original equation

OpenStudy (yadu123):

or you can add one to the other side to make it positive but thats an extra step

OpenStudy (kitkat1):

Okay so 4x - 9 = -1 /4 /4 Right?

OpenStudy (yadu123):

no first you add 9. 9+-1=8

OpenStudy (kitkat1):

What did u add?\

OpenStudy (yadu123):

So you understand how we got 4x-9=-1 Right?

OpenStudy (kitkat1):

yes but what is next

OpenStudy (yadu123):

So to solve for x bring 9 to the other side. Since 9 is negative on the other side you do the opposite and add

OpenStudy (yadu123):

4x=9+-1

OpenStudy (yadu123):

9+-1 is 8. Divide by 4 and get 2

OpenStudy (kitkat1):

So the solution is 2

OpenStudy (yadu123):

x=2

OpenStudy (kitkat1):

okay can u help me with one more?

OpenStudy (yadu123):

sure

OpenStudy (kitkat1):

A pair of equations is shown below: y = 7x – 9 y = 3x – 1 Part A: In your own words, explain how you can solve the pair of equations graphically. Write the slope and y-intercept for each equation that you will plot on the graph to solve the equations.

OpenStudy (yadu123):

Ok same thing here set each equation equal to each other

OpenStudy (kitkat1):

so 7x-9 = 3x - 1?

OpenStudy (yadu123):

Yes. Now combine like terms

OpenStudy (kitkat1):

do I add the 7 or 3

OpenStudy (yadu123):